# Bridging Domain Gap for Flight-Ready Spaceborne Vision

### v1

@@ Abstract
This work presents Spacecraft Pose Network v3 (SPNv3), a Neural Network (NN) for monocular pose estimation of a known, non-cooperative target spacecraft. As opposed to existing literature, SPNv3 is designed and trained to be computationally efficient while providing robustness to spaceborne images that have not been observed during offline training and validation on the ground. These characteristics are essential to deploying NNs on space-grade edge devices. They are achieved through careful NN design choices, and an extensive trade-off analysis reveals features such as data augmentation, transfer learning and vision transformer architecture as a few of those that contribute to simultaneously maximizing robustness and minimizing computational overhead. Experiments demonstrate that the final SPNv3 can achieve state-of-the-art pose accuracy on hardware-in-the-loop images from a robotic testbed while having trained exclusively on computer-generated synthetic images, effectively bridging the domain gap between synthetic and real imagery. At the same time, SPNv3 runs well above the update frequency of modern satellite navigation filters when tested on a representative graphical processing unit system with flight heritage. Overall, SPNv3 is an efficient, flight-ready NN model readily applicable to a wide range of close-range rendezvous and proximity operations with target resident space objects. The code implementation of SPNv3 will be made publicly available.
@@ 1 Introduction
The autonomous Rendezvous, Proximity Operations and Docking (RPOD) capability with non-cooperative Resident Space Objects (RSO) is a core technological requirement for various future space missions. Aimed at sustainable space development, these missions include on-orbit servicing and refueling operations such as now-canceled OSAM missions by NASA [1, 2] and active debris removal such as RemoveDEBRIS [3] by Surrey Space Center, ADRAS-J [4] by Astroscale and ClearSpace-1 [5] by ClearSpace SA. The RPOD capability with such targets—defunct satellites, debris, etc.—requires accurate, real-time knowledge of the position and orientation (i.e., pose) of the target spacecraft with respect to the servicer spacecraft. In these scenarios, the non-cooperative nature of the target implies that it is without active communication links or aiding fiduciary markers on its surface to guide the rendezvous process, such that the servicer must be able to estimate and track the target’s pose using the onboard sensors and computers only. Given such constraints, a monocular camera is an attractive choice of sensor due to its low Size-Weight-and-Power-Cost (SWaP-C), which is suitable for miniaturized systems such as SmallSats and CubeSats. Moreover, low SWaP-C implies additional sensor redundancy and fail-proofing compared to more complex sensor systems such as LIght Detection And Ranging (LIDAR) and stereovision which requires careful calibration.
@@
Recent years have seen a significant breakthrough in Machine Learning (ML)-based monocular pose estimation of a known, non-cooperative spacecraft. Training of these ML models such as Convolutional Neural Networks (CNN) generally requires a large-scale dataset of images and pose labels. However, access to space is difficult and expensive, and a close-range rendezvous of two satellites is in itself a very rare occasion, making in-situ data collection and label annotation a challenging task for spaceborne applications. In response, the aerospace community has adopted a methodology whereby synthetic images are generated with rendering tools and gaming engines such as OpenGL and Unreal Engine in order to train the neural networks [6, 7, 8, 9]. However, synthetic images have inherently dissimilar visual features compared to real spaceborne images characterized by low Signal-to-Noise Ratio (SNR) and harsh illumination conditions. In addition, the degradation of vision cameras and target RSOs due to operations in the space environment is difficult to model prior to the mission and through computer graphics. This results in the problem known as domain gap, which manifests as a drop in performance when the models are tested on the Out-Of-Distribution (OOD) data sampled from a statistical distribution different from that of the training data [10, 11]. In case the models are trained with synthetic images and tested on real-life images, it is also known as reality gap [12] or sim2real gap [13] primarily in robotics literature.
@@
The inaccessibility of space poses a particularly challenging logistics problem to not only overcoming the domain gap but also verifying it during the stringent pre-flight verification processes typically required of space missions. The most prominent approach taken by the aerospace community is to utilize a robotic testbed to simulate the space environment on-ground. The idea is to physically stimulate the vision-based sensors using a mockup model of the target placed in a facility simulating the high-fidelity space-like illumination conditions. A well-calibrated testbed can then be used to re-create various RPOD scenarios, generating pose-annotated images at scale with minimal human intervention for the purpose of validating a NN’s robustness across the sim2real gap. One such example is the Testbed for Rendezvous and Optical Navigation (TRON) facility at the Stanford’s Space Rendezvous Laboratory (SLAB) [^3] https://slab.stanford.edu/ [14] which was used to create the so-called Hardware-In-the-Loop (HIL) images that constitute datasets such as SPEED+ [15] and SHIRT [16]. Figure 1 shows the TRON facility and a few example images from the two HIL domains of the SPEED+ dataset—lightbox and sunlamp. In contrast to SPEED+ which contains static images at random poses, the SHIRT dataset includes sequential lightbox images obtained during dynamic RPOD scenarios. This allows the testing of navigation filters with ML algorithms in the loop. Overall, these HIL images can then be used as on-ground surrogates of otherwise unavailable spaceborne images for the evaluation of NN and the navigation algorithm robustness across domain gaps. SPEED+ was used as the core dataset of the second Satellite Pose Estimation Competition (SPEC2021) [^4] https://kelvins.esa.int/pose-estimation-2021/ co-organized by SLAB and the Advanced Concepts Team (ACT) of the European Space Agency (ESA), in which the participants were tasked to predict the poses on SPEED+ HIL images while only having access to the labeled synthetic and unlabeled HIL domain images [17].
@@
Given such tools and datasets, the authors have explored a strategy to fully close the domain gap between synthetic and spaceborne images. The strategy consists of three steps. The first is to train a NN model using only synthetic images to be as robust as possible to the unknown spaceborne images [18]. The robustness is evaluated on-ground using HIL domain images which are excluded from the training. Then, the NN is integrated as a measurement module into an Adaptive Unscented Kalman Filter (AUKF) which tracks the pose of the target [16]. While it was shown that a well-designed AUKF allows robust tracking of the target’s pose despite the domain gap suffered by its NN measurement module, there is still a remaining gap between the HIL imagery used for on-ground evaluation and the spaceborne flight images. The primary reason is that HIL images use an inexpensive mockup model of the target spacecraft which has different material and surface properties compared to the real satellite. Therefore, Park and D’Amico [19] proposed to additionally fine-tune the NN weights using the flight images collected during in-space RPOD. The fine-tuning is done online in a supervised manner whereby the pose pseudo-labels are obtained from the most up-to-date state estimates of the onboard AUKF. This work has shown that Online Supervised Training (OST) can fully close the domain gap in the orientation prediction even when using a sub-optimally trained NN.
Despite the success of OST, there are benefits to further maximizing the robustness of onboard NN during the offline training on synthetic images. The first is that OST requires that the onboard ML-in-the-loop navigation filter be able to first converge to steady-states so that accurate pose pseudo-labels can be generated from the state estimates. The probability of failure to converge is minimized as the OOD robustness of the onboard NN is maximized prior to deployment. The second is that training a NN onboard the limited computing environment of satellite avionics could be expensive both in terms of the computational efficiency and power consumption. This could pose a logistical challenge to schedule the NN training amidst the nominal Guidance, Navigation and Control (GN&C) operations. These two reasons motivate
@@
    a training mechanism that renders the NN as robust as possible across domain gap from the offline training alone, which effectively minimizes the required number of OST steps performed in space, and
@@
These two objectives—OOD robustness and computational efficiency—are at odd, since it is well known that larger NNs with higher capacity tend to perform better on both in-distribution and OOD data [20, 21].
In response to the above challenge, the main contribution of this work is Spacecraft Pose Network v3 (SPNv3), a NN model based on Vision Transformer (ViT) [22] for monocular pose estimation of a non-cooperative spacecraft. SPNv3 is designed and trained with onboard computational efficiency and robustness across the sim2real gap as top priorities. Extensive ablation studies of SPNv3 on the SPEED+ dataset demonstrate that ViT-based models without convolution operations dominate the computational efficiency front, and that their OOD robustness on HIL domain images can be improved by employing extensive data augmentation, enhanced transfer learning and increased input image resolution with minimal gain of computational overhead for inference. A comprehensive set of experiments reveals that the proposed SPNv3 can achieve state-of-the-art performances on the HIL domain images of the SPEED+ dataset while training exclusively on its synthetic images. Most importantly, a mid-size variant of SPNv3 would rank the first place on the lightbox category of SPEC2021 without any adversarial training or unsupervised domain adaptation, the methods which all winners employed by directly including the unlabeled test domain images into the training process. Such state-of-the-art robustness is achieved while taking no more than 40 ms per inference on an NVIDIA Jetson Nano 4GB [23], a representative, restricted compute environment similar to GPU-powered systems on satellite missions [24, 25].
This paper is organized as follows. Section 2 provides a brief overview of literature on ML-based monocular pose estimation and existing methods to bridge domain gap. Section 3 then goes over three key requirements that must be met by spaceborne ML models in addition to OOD robustness, two of which drive various design and training options for SPNv3 introduced in section 4. Section 5 shows extensive experiments on the SPEED+ dataset to identify key elements of the design and training of SPNv3 that contribute the most to its robustness on HIL domain images and computational efficiency. The paper ends with conclusions and future works in Section 6.
@@ 2.1 ML Approaches to Spacecraft Pose Estimation
The first ML-based approach to pose estimation of a known target spacecraft was Spacecraft Pose Network (SPN) [6] which performs 1) relative attitude determination via a hybrid approach of attitude classification and regression and 2) translation estimation by exploiting the perspective transformation and geometric constraints. Since it is impossible to collect a pose-annotated image dataset from space, the same work introduced the Spacecraft PosE Estimation Dataset (SPEED) [26] which consists of 15,300 images of the Tango spacecraft from the PRISMA mission [27]. Specifically, the dataset comprises 1) 15,000 synthetic images rendered with OpenGL-based Optical Stimulator (OS) camera emulator software [28, 29] of SLAB multi-Satellite Software Simulator ($S^{3}$) [30] and 2) 300 real images of a mockup of the same target captured from the TRON facility at SLAB. The dataset was made publicly available as part of the first Satellite Pose Estimation Challenge (SPEC2019) [^5] https://kelvins.esa.int/satellite-pose-estimation-challenge/ [31] co-hosted by SLAB and ESA.
@@
[…]
In addition to SPEED, a number of datasets for spacecraft pose estimation have been published and made publicly available in the literature. For example, Proença and Gao [7], who also ranked third place in SPEC2019, published the Unreal Rendered Spacecraft On-orbit (URSO) dataset which uses Unreal Engine 4 to render synthetic images of the Soyuz and Dragon spacecraft. Kaki et al. [37] renders synthetic images of the Cygnus spacecraft using Blender and its Cycles rendering engine. Dung et al. [40] also renders about 3,000 synthetic images of different satellites for various computer vision tasks such as bounding box prediction and satellite foreground and parts segmentation. Other authors also created their own synthetic datasets, such as those of the Envisat spacecraft rendered with Cinema 4D [38], the SwissCube dataset for spacecraft pose estimation from wide-depth-range images using the Mitsuba 2 renderer [36], the SPARK [8] dataset which contains images of 11 different spacecraft rendered with the Unity3D game engine, and Synthetic-Minerva II2 [41] which renders images of the Minerva-II2 rover from the Hayabusa2 mission [42] using SolidWorks’ Photoview 360 renderer.
@@ 2.2 Algorithms to Bridge Domain Gap
The general approaches to tackling domain gaps can be put into two categories: domain adaptation and domain randomization. Specifically, Unsupervised Domain Adaptation (UDA) aims to close the gap by directly incorporating the unlabeled target domain images into the training phase [10, 43, 44, 45, 11, 46, 47, 48, 49]. Many different approaches have been proposed in the literature to close the said gap. One primary approach is to align the source and the target data distributions in the latent feature space by minimizing a certain divergence criterion representing the domain discrepancy, such as maximum mean discrepancy [50, 45, 51, 49], $\mathcal{H}$-divergence [10, 44], correlation via statistical moments [52, 46] or Wasserstein distance [53, 54, 55]. Another approach employs adversarial training [44, 56, 48] so that the predicted features become indistinguishable for both source and target domain inputs from the perspective of an auxiliary domain discriminator network. Note that the aforementioned approaches all aim to promote learning invariant representations by aligning both source and target domain data in the feature space.
One major shortcoming of UDA is that it requires simultaneous availability of the labeled source and unlabeled target domain data. This assumption violates the operational constraints of space missions as spaceborne images of the target do not become available until on-orbit rendezvous. Moreover, the processor and memory onboard the satellites are extremely limited for the full training session using large-scale synthetic source images and spaceborne target images collected during RPOD. Therefore, the conventional UDA approaches cannot be used for realistic space mission scenarios.
As opposed to UDA, domain randomization instead aims to randomize various aspects of the training images such that OOD test images would appear as another randomized instance of the training set [57, 58, 59, 60, 61]. For example, Sadeghi and Levine [58] and Tobin et al. [59] randomize lighting conditions, object textures and placements at rendering stage to train domain randomized reinforcement learning models. Jackson et al. [62] proposes style augmentation which randomizes the image texture [63] via neural style transfer [64]. While domain randomization does not require target domain images during the training phase, it also makes it difficult to provide any assurance that the CNN trained with domain randomization will work well on the target spaceborne images.

### v2

# Bridging the Domain Gap for Flight-Ready Spaceborne Vision

@@ Abstract
This work presents Spacecraft Pose Network v3 (SPNv3), a Neural Network (NN) for monocular pose estimation of a known, non-cooperative target spacecraft. SPNv3 is designed and trained to be computationally efficient while providing robustness to spaceborne images that have not been observed during offline training and validation on the ground. These characteristics are essential to deploying NNs on space-grade edge devices. They are achieved through careful NN design choices, and an extensive trade-off analysis reveals features such as data augmentation, transfer learning and vision transformer architecture as a few of those that contribute to simultaneously maximizing robustness and minimizing computational overhead. Experiments demonstrate that the final SPNv3 can achieve state-of-the-art pose accuracy on hardware-in-the-loop images from a robotic testbed while having trained exclusively on computer-generated synthetic images, effectively bridging the domain gap between synthetic and real imagery. At the same time, SPNv3 runs well above the update frequency of modern satellite navigation filters when tested on a representative graphical processing unit system with flight heritage. Overall, SPNv3 is an efficient, flight-ready NN model readily applicable to close-range rendezvous and proximity operations with target resident space objects.
22footnotetext: This work was done as part of the Ph.D. program of T. H. Park at the Space Rendezvous Laboratory (SLAB), Stanford University.
@@ 1 Introduction
Autonomous Rendezvous, Proximity Operations and Docking (RPOD) capability with non-cooperative Resident Space Objects (RSO) is a core technological requirement for various future space missions. Aimed at sustainable space development, these missions include on-orbit servicing and refueling operations such as the now-canceled OSAM missions by NASA [1, 2] and active debris removal such as RemoveDEBRIS [3] by Surrey Space Center, ADRAS-J [4] by Astroscale and ClearSpace-1 [5] by ClearSpace SA. The RPOD capability with such targets—defunct satellites, debris, etc.—requires accurate, real-time knowledge of the position and orientation (i.e., pose) of the target spacecraft with respect to the servicer spacecraft. In these scenarios, the non-cooperative nature of the target implies that it is without active communication links or aiding fiduciary markers on its surface to guide the rendezvous process, such that the servicer must be able to estimate and track the target’s pose using the onboard sensors and computers only. Given such constraints, a monocular camera is an attractive choice of sensor due to its low Size-Weight-and-Power-Cost (SWaP-C), which is suitable for miniaturized systems such as SmallSats and CubeSats. Moreover, low SWaP-C implies additional sensor redundancy and fail-proofing compared to more complex sensor systems such as LIght Detection And Ranging (LIDAR) and stereovision. In particular, LIDAR allows accurate range detection at far range [6, 7], but it is subject to much stricter SWaP-C requirements. On the other hand, while stereovision utilizes low SWaP-C vision sensors amenable to SmallSats [8, 9], its applicability for range estimation beyond proximity is restricted by the requirements for a large baseline, synchronized measurements and careful calibration.
@@
Recent years have seen a significant breakthrough in Machine Learning (ML)-based monocular pose estimation of a known, non-cooperative spacecraft. Training of these ML models such as Convolutional Neural Networks (CNN) generally requires a large-scale dataset of images and pose labels. However, access to space is difficult and expensive, and a close-range rendezvous of two satellites is, in itself, a very rare occasion, making in-situ data collection and label annotation a challenging task for spaceborne applications. In response, the aerospace community has adopted a methodology whereby synthetic images are generated with rendering tools and gaming engines such as OpenGL and Unreal Engine in order to train the neural networks [10, 11, 12, 13]. However, synthetic images have inherently dissimilar visual features compared to real spaceborne images characterized by low Signal-to-Noise Ratio (SNR) and harsh illumination conditions. In addition, the degradation of vision cameras and target RSOs due to operations in the space environment is difficult to model prior to the mission and through computer graphics. This results in the problem known as domain gap, which manifests as a drop in performance when the models are tested on the Out-Of-Distribution (OOD) data sampled from a statistical distribution different from that of the training data [14, 15]. In case the models are trained with synthetic images and tested on real-life images, it is also known as reality gap [16] or sim2real gap [17] primarily in the robotics literature.
@@
The inaccessibility of space poses a particularly challenging logistics problem to not only overcoming the domain gap but also verifying it during the stringent pre-flight verification processes typically required of space missions. The most prominent approach taken by the aerospace community is to utilize a robotic testbed to simulate the space environment on ground. The idea is to physically stimulate the vision-based sensors using a mockup model of the target placed in a facility simulating the high-fidelity space-like illumination conditions. A well-calibrated testbed can then be used to re-create various RPOD scenarios, generating pose-annotated images at scale with minimal human intervention for the purpose of validating a NN’s robustness across the sim2real gap. One such example is the Testbed for Rendezvous and Optical Navigation (TRON) facility at Stanford’s Space Rendezvous Laboratory (SLAB) [^3] https://slab.stanford.edu/ [18], which was used to create the so-called Hardware-In-the-Loop (HIL) images that constitute datasets such as SPEED+ [19] and SHIRT [20]. Figure˜1 shows the TRON facility and a few example images from the two HIL domains of the SPEED+ dataset—lightbox and sunlamp. In contrast to SPEED+, which contains static images at random poses, the SHIRT dataset includes sequential lightbox images obtained during dynamic RPOD scenarios. This allows the testing of navigation filters with ML algorithms in the loop. Overall, these HIL images can then be used as on-ground surrogates of otherwise unavailable spaceborne images for the evaluation of NN and the navigation algorithm robustness across domain gaps.
SPEED+ was used as the core dataset for the second Satellite Pose Estimation Competition (SPEC2021) [^4] https://kelvins.esa.int/pose-estimation-2021/ co-organized by SLAB and the Advanced Concepts Team (ACT) of the European Space Agency (ESA). From October 2021 to March 2022, participants were tasked to predict poses on SPEED+ HIL images while having access only to the labeled synthetic and unlabeled HIL domain images [21]. The submitted pose predictions were then evaluated based on a combination of translation and orientation errors separately for the lightbox and sunlamp domains. Finally, after the conclusion of the official competition, a separate post-mortem challenge was held until the end of 2022 in order to attract more innovative solutions for bridging the domain gap in spaceborne imagery.
@@
Given such tools and datasets, the authors have explored a strategy to fully close the domain gap between synthetic and spaceborne images. The strategy consists of three steps. The first is to train a NN model using only synthetic images to be as robust as possible to the unknown spaceborne images [22]. The robustness is evaluated on ground using HIL domain images that are excluded from the training. Then, the NN is integrated as a measurement module into an Adaptive Unscented Kalman Filter (AUKF), which tracks the pose of the target [20]. While it was shown that a well-designed AUKF allows robust tracking of the target’s pose despite the domain gap suffered by its NN measurement model, there is still a remaining gap between the HIL imagery used for on-ground evaluation and the spaceborne flight images. The primary reason is that HIL images use an inexpensive mockup model of the target spacecraft which has different material and surface properties compared to the real satellite. Therefore, Park and D’Amico [23] proposed to additionally fine-tune the NN weights using the flight images collected during in-space RPOD. The fine-tuning is done online in a supervised manner whereby the pose pseudo-labels are obtained from the most up-to-date state estimates of the onboard AUKF. This work has shown that Online Supervised Training (OST) can fully close the domain gap in the orientation prediction even when using a sub-optimally trained NN.
Despite the success of OST, there are benefits to further maximizing the robustness of an onboard NN during the offline training on synthetic images. The first is that OST requires that the onboard ML-in-the-loop navigation filter be able to first converge to steady state so that accurate pose pseudo-labels can be generated from the state estimates. The probability of failure to converge is minimized as the OOD robustness of the onboard NN is maximized prior to deployment. The second is that training a NN onboard the limited computing environment of satellite avionics could be expensive both in terms of the computational effort and power consumption. This could pose a logistical challenge to schedule the NN training amidst the nominal Guidance, Navigation and Control (GN&C) operations. These two reasons motivate
@@
    a training mechanism that renders the NN as robust as possible across the domain gap from the offline training alone, which effectively minimizes the required number of OST steps performed in space, and
@@
These two objectives—OOD robustness and computational efficiency—are at odds, since it is well known that larger NNs with higher capacity tend to perform better on both in-distribution and OOD data [24, 25].
In response to the above challenge, the main contribution of this work is Spacecraft Pose Network v3 (SPNv3), a NN model based on Vision Transformer (ViT) [26] for monocular pose estimation of a non-cooperative spacecraft. SPNv3 is designed and trained with onboard computational efficiency and robustness across the sim2real gap as top priorities. Extensive ablation studies of SPNv3 on the SPEED+ dataset demonstrate that ViT-based models without convolution operations dominate on the computational efficiency front, and that their OOD robustness on HIL domain images can be improved by employing extensive data augmentation, enhanced transfer learning and increased input image resolution with minimal gain of computational overhead for inference. A comprehensive set of experiments reveals that the proposed SPNv3 can achieve state-of-the-art performances on the HIL domain images of the SPEED+ dataset while training exclusively on its synthetic images. Most importantly, a mid-size variant of SPNv3 would rank the first place on the lightbox category of SPEC2021 without any adversarial training or unsupervised domain adaptation, the methods which all winners employed by directly including the unlabeled test domain images into the training process. Such state-of-the-art robustness is achieved while taking no more than 40 ms per inference on an NVIDIA Jetson Nano 4GB [27], a representative, restricted compute environment similar to GPU-powered systems on satellite missions [28, 29].
This paper is organized as follows: Section˜2 provides a brief overview of literature on ML-based monocular pose estimation and existing methods to bridge domain gap. Section˜3 then goes over three key requirements that must be met by spaceborne ML models in addition to OOD robustness. Two of them drive various design and training options for SPNv3 introduced in section˜5 after briefly explaining the adopted pose estimation pipeline and assumptions in section˜4. Section˜6 shows extensive experiments on the SPEED+ dataset to identify key elements of the design and training of SPNv3 that contribute the most to its robustness on HIL domain images and computational efficiency. The paper ends with conclusions and future works in Section˜7.
@@ 2.1 ML Approaches to Spacecraft Pose Estimation
One of the first ML-based approaches to pose estimation of a known target spacecraft was Spacecraft Pose Network (SPN) [30, 31, 10] which performs 1) relative attitude determination via a hybrid approach of attitude classification and regression and 2) translation estimation by exploiting the perspective transformation and geometric constraints. Since it is impractical to collect a pose-annotated image dataset from space, the same work introduced the Spacecraft PosE Estimation Dataset (SPEED) [32] which consists of 15,300 images of the Tango spacecraft from the PRISMA mission [33]. Specifically, the dataset comprises 1) 15,000 synthetic images rendered with OpenGL-based Optical Stimulator (OS) camera emulator software [34, 31] of SLAB multi-Satellite Software Simulator ($S^{3}$) [35] and 2) 300 real images of a mockup of the same target captured from the TRON facility at SLAB. The dataset was made publicly available as part of the first Satellite Pose Estimation Challenge (SPEC2019) [^5] https://kelvins.esa.int/satellite-pose-estimation-challenge/ [36] co-hosted by SLAB and ESA.
@@
Many top-performing entries of SPEC2019 have adopted diverse CNN architectures and pose estimation strategies, such as probabilistic orientation estimation via soft classification [11] or estimation of 2D pixel coordinates of the designated keypoints on the spacecraft surface [37, 38]. Specifically, Chen et al. [38] and Park et al. [37], who respectively ranked first and fourth places in the challenge, independently proposed a three-stage architecture: 1) an object detection CNN which is used to identify the Region-of-Interest (RoI) around the target, 2) a pose estimation CNN which takes in an image cropped around the detected RoI and outputs the 2D keypoint locations, and 3) a PnnP module which solves for the full 6D pose based on the known correspondence of detected keypoint 2D locations and 3D model coordinates. Notably, Park et al. [37] directly regresses $(x,y)$ coordinates of the keypoints, whereas Chen et al. [38] outputs a set of 2D heatmaps whose peaks correspond to the locations of the keypoints. They also showed that cropping the input image around the target using a detected bounding box is crucial to performing pose estimation of a far-away target. For example, the original resolution of the SPEED images is 1920 $\times$ 1200, while most CNNs for ImageNet [39] classification expect 224 $\times$ 224 inputs. Cropping around the target prior to such dramatic down-scaling helps preserve much of the detailed target features that would otherwise be lost due to large inter-spacecraft separation as visualized in Fig.˜2. Many CNN models that followed SPEC2019 also adopt similar strategies [40, 41, 42, 12, 43, 44]. Readers are referred to Pasqualetto Cassinis et al. [45] for a more comprehensive review of monocular spacecraft pose estimation using both conventional and deep learning-based methods.
In addition to SPEED, a number of datasets for spacecraft pose estimation have been published and made publicly available in the literature. For example, Proença and Gao [11], who also ranked third place in SPEC2019, published the Unreal Rendered Spacecraft On-orbit (URSO) dataset, which uses Unreal Engine 4 to render synthetic images of the Soyuz and Dragon spacecraft. Kaki et al. [42] renders synthetic images of the Cygnus spacecraft using Blender and its Cycles rendering engine. Dung et al. [46] also renders about 3,000 synthetic images of different satellites for various computer vision tasks such as bounding box prediction and satellite foreground and parts segmentation. Other authors also created their own synthetic datasets, such as those of the Envisat spacecraft rendered with Cinema 4D [43], the SwissCube dataset for spacecraft pose estimation from wide-depth-range images using the Mitsuba 2 renderer [41], the SPARK [12] dataset which contains images of 11 different spacecraft rendered with the Unity3D game engine, and Synthetic-Minerva II2 [47] which renders images of the Minerva-II2 rover from the Hayabusa2 mission [48] using SolidWorks’ Photoview 360 renderer.
@@ 2.2 Algorithms to Bridge Domain Gap
Approaches to tackling domain gaps can be put into two categories: domain adaptation and domain randomization. Specifically, Unsupervised Domain Adaptation (UDA) methods aim to close the gap by directly incorporating the unlabeled target-domain images into the training phase [14, 49, 50, 51, 15, 52, 53, 54, 55]. One primary approach is to align the source and the target data distributions in the latent feature space by minimizing a certain divergence criterion representing the domain discrepancy, such as maximum mean discrepancy [56, 51, 57, 55], $\mathcal{H}$-divergence [14, 50], correlation via statistical moments [58, 52] or Wasserstein distance [59, 60, 61]. Another approach employs adversarial training [50, 62, 54] so that the predicted features become indistinguishable for both source and target domain inputs from the perspective of an auxiliary domain discriminator network. Note that the aforementioned approaches all aim to promote learning invariant representations by aligning both source and target domain data in the feature space.
One major shortcoming of UDA is that it requires simultaneous availability of the labeled source and unlabeled target-domain data. This assumption violates the operational constraints of space missions, as spaceborne images of the target do not become available until on-orbit rendezvous. Moreover, the processor and memory onboard the satellites are extremely limited for the full training session using large-scale synthetic source images and spaceborne target images collected during RPOD. Therefore, the conventional UDA approaches cannot be used for realistic space mission scenarios.
As opposed to UDA, domain randomization instead aims to randomize various aspects of the training images such that OOD test images would appear as another randomized instance of the training set [63, 64, 65, 66, 67]. For example, Sadeghi and Levine [64] and Tobin et al. [65] randomize lighting conditions, object textures and placements at the rendering stage to train domain randomized reinforcement learning models. Jackson et al. [68] proposes style augmentation, which randomizes the image texture [69] via neural style transfer [70]. While domain randomization does not require target domain images during the training phase, it also makes it difficult to provide any assurance that the CNN trained with domain randomization will work well on the target spaceborne images.